OpenStudy (anonymous):

A 18 gram sample of a substance that's used to detect explosives has a k-value of 0.125 find the substances half-life, in days. Round your answers to the nearest tenth

OpenStudy (anonymous):

T1/2 = ln 2 / k = ln 2 / 0.125 = ???????

OpenStudy (anonymous):

=5.5?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

okay thanks